OpenStudy (anonymous):

Sally wants to buy some accessories for her hair. She can buy two scrunchies and four barrettes for $12.50 or one scrunchie and six barrettes for $12.25. How much does the scrunchie cost that she wants to buy?

OpenStudy (anonymous):

x-scrunchies prize y-barrettes prize 2x+4y=12,50 1x+6y=12,25 solve this system

OpenStudy (apoorvk):

Let the cost of one scrunchie be 'x' and that of one barrete be 'y'. So, cost of 2 scrunchies and 4 barretes = 2x + 4y = 12.5 ...{i} and, cost of 1 scrunchie and 6 barretes = x + 6y =12.25 ...{ii} Solve this system^, and find 'x'!

OpenStudy (apoorvk):

P.S. - I have no idea what scrunchies or barretes are! -.- Sounds more like food to me :P

OpenStudy (anonymous):

There for hair

OpenStudy (apoorvk):

haha, yeah. So understanding this? getting an answer?

OpenStudy (anonymous):

No